OpenStudy (anonymous):

Decompose into partial fractions: (x^3+2)/(x^2-1)

OpenStudy (anonymous):

\[\LARGE \frac{x^3+2}{x^2-1}=\frac{x^3+2}{(x-1)(x+1)}=\] \[\LARGE {x^3+2\over x-1}\cdot {1\over x+1}=\left( {x^3\over x-1}+{2\over x-1 }\right)\cdot {1\over x+1} \] I don't know if this is what you're asking for, if it's not, throw it away ;)

OpenStudy (anonymous):

divide first

OpenStudy (anonymous):

Its fine to the point\[(x^3+2)/(x-1)(x+1)\] but now I have to solve for \[(x^3+2)/(x-1)(x+1) = A/(x-1) + B/(x+1)\] want to check ifthe answers I got for A and B are correct

OpenStudy (anonymous):

divide first to get \[x+\frac{x+2}{(x+1)(x-1)}\]

OpenStudy (anonymous):

the degree of the numerator is bigger than the degree of the denominator,so before trying partial fraction decomposition divide to get a quotient and a remainder. then pfd for the remainder

OpenStudy (anonymous):

what experimentx said before he/she deleted it

OpenStudy (anonymous):

\[\frac{x^3+2}{x^2-1}=x+\frac{3}{2 (x-1)}-\frac{1}{2 (x+1)} \]

OpenStudy (anonymous):

Thanks that helped alot

OpenStudy (anonymous):

@satellite73 I will always remember your words before I attempt a partial fraction question again

OpenStudy (anonymous):

glad to help